Support Services Meeting Notes
June 11, 2007
June 11, 2007
Present: Bill Covey, Kelley Cunningham, Barbara Oliver, Amy Polk, Rob Roberts, Laurie TaylorBusiness Services
- Salary budget is in - it balanced perfectly.
- POs have closed. Any POs trying to go through after 5PM on Friday - you have to prove that it is a true emergency or it will be cancelled within two days.
- New DSO (Direct Service Organization) procedures - it has been decided that all independent financial groups will be put on PeopleSoft. Some procedures that we use will require re-training.
- We are trying to get all bills paid.
- Orders will not open again until July 10.
- VOIP in East - no date yet.
- Waiting on final year things for room 450 East (Special Collections) - carpet, construction, shelving.
- Men's room sign for downstairs (West) - they sent the wrong one (the old signage color). We will not be charged. We've made the correction.
- Carrels - 4th floor is done with the exception of one adjustment on a door. Only have a few minor adjustments to make on the 2nd floor. All the locks are completely done. Rob should send an email to Lori (copying Bill) when the carrels are completely finished.
- Space allocation is done.
- Asset inventory is done.
- Insurance replacement survey - finished. Will get it to Barbara Oliver this week.
- UPS new account for ILL - won't let them put it on the PCard.
- Chairs for the lobby are in - will distribute. The old ones need cleaning. Need to get an estimate.
- Wood You shelving was delivered. It is sagging already. They are solid oak shelves. The company is going to come out and take a look at it. Support Services is to pick out all furniture from now on.
- This will be the second year of the mini-grant program. We are updating the procedures (re-evaluating the process). The deadline last year was the end of September. We'll get an announcement out. Barbara will have to figure out how to get the money out of DSO.
- LSTA - Erich's grant - did it get sent back? The word was given. Don't know if it has been done. Need to check with Grace.
- All orders have been submitted.
- The new replacement file and print server complex was ordered. It will have live replication in two places. Should be much faster. There will be some tape activity, but not as much. Heavy virtualization.
- The addition of memory to the Web server seems to have taken care of the DLC Web problems. They have all been working fine. No time differences in performance. Occasional slowness, but no outages.
- Fire in Education - a fan in the computer lab downstairs. Smoke distributed throughout the building. HVAC contacted. The smell was like polyvinyl burning. The sprinklers came on. We got the smoke out. Education also said the power was out. Bill checked again on Sunday evening. Preservation will check out the collection for damage in a day or two. Most of the damage was downstairs (not us). The library was re-opened today.
- Directors - have not been in the office much of late - only one or two of us at a time. Judy has been on the road, John has been traveling and Carol has been on vacation. We will be looking at the budget for next year: ways to pay for computers and expand the security guard service. We want to expand the guard to an eight-hour shift to include the branch libraries.
- ALA is coming up - librarians will be out at the conference.